OpenStudy (anonymous):

What would be the mean value of the function sin x over the range of values x=0 to x=pi ?

OpenStudy (anonymous):

Hmm the answer is $$ \frac 2 \pi\large\int \limits_{0} ^\frac \pi 2 \sin x $$ can you do the rest ?